we keep the switch at position x for a sufficient time and thereafter (at t0) connect it to Y, in the circuit shown below.






what amount of current will be measured through the resistor after 5.0 micro-second from t0, if V=9v, R=10 and C=1


1
Expert's answer
2021-08-20T12:55:27-0400

The voltage across the capacitor will be


"V_C=V_s\big(1-e^{-t/(RC)}\big)."

The current:


"I=\frac{V_C}{R}=\frac{V_s\big(1-e^{-t/(RC)}\big)}{R}=0.35\text{ A}."